E7.20 (LO 5) (Analysis of Receivables) Presented below is information for Jones Company. 1. Beginning-of-the-year Accounts Receivable balance was $15,000. 2. Net sales (all on account) for the year were $100,000. Jones does not offer cash discounts. 3. Collections on accounts receivable during the year were $70,000. Instructions a. Prepare (summary) journal entries to record the items noted above. b. Compute Jones’s accounts receivable turnover and days to collect receivables for the year. The company does not believe it will have any bad debts. c. Use the turnover ratio computed in (b) to analyze Jones’s liquidity. The turnover ratio last year was 6.0.
Answer #1

a.

JOURNAL
Date General Journal Debit Credit
1 No journal entry needed
No journal entry needed
2 Accounts receivable 100,000
Sales 100,000
3 Cash 70,000
Accounts receivable 70,000

b.

Ending accounts receivable = Beginning accounts receivable + Sales - Cash received from accounts receivable

= 15,000+100,000-70,000

= $45,000

Average accounts receivable = ( Beginning accounts receivable +Ending accounts receivable )/2

= (15,000+45,000)/2

= $30,000

Accounts receivable turnover ratio = Sales /Average accounts receivable

= 100,000/30,000

= 3.33 times

Days to collect = 365/ Accounts receivable turnover ratio

= 365/3.33

= 110 days

c.

In the current year, liquidity has decreased since accounts receivable turnover ratio has decreased in current year.

Kindly comment if you need further assistance. Thanks‼!